# Applying a STEM Ways of Thinking Framework for Student-generated Engineering Design-based Physics Problems

### v1

@@
This second paper in a multi-part series builds on the first, which introduced the Ways of Thinking for Engineering Design-based Physics (WoT4EDP) framework for STEM education in an introductory undergraduate physics course. Here, we apply the framework to analyze transcripts of group discussions and written reports from 14 student teams as they engaged in a self-generated engineering design (ED) problem in an introductory physics laboratory. We qualitatively examine: (i) the aspects students address in their problem statements; (ii) how they engage in design-based, science-based, and mathematics-based thinking, as well as metacognitive reflection, while developing solutions; and (iii) how they incorporate computational thinking through Python coding. Key findings highlight the need for: (i) increased guidance for iterative problem framing; (ii) structured support for assessing design limitations, engaging in a feasibility study, adopting a systematic approach to applying physics and mathematics in their iterations, and making specific metacognitive reflections; and (iii) integration of Python-based activities into laboratory tasks with appropriate scaffolding.
We present our findings through a detailed qualitative analysis, drawing extensively from the qualitative methods literature. We outline our analytical approach, present coding charts, and employ qualitative methods such as thematic analysis and thick description to convey our findings. In doing so, we contribute to ongoing efforts to enhance the rigor of qualitative analysis in physics education research (PER). Based on our analysis, we provide valuable insight for educators and researchers in designing physics-based engineering design tasks and promoting interdisciplinary problem-solving in STEM education.
@@
Ravishankar Chatta Subramaniam Department of Physics and Astronomy, Purdue University, 525 Northwestern Ave, West Lafayette, IN 47907, U.S.A.    Nikhil Borse Dept. of Physics and Astronomy, Purdue University, West Lafayette, IN 47907, U.S.A.    Winter Allen Dept. of Physics and Astronomy, Purdue University, West Lafayette, IN 47907, U.S.A.    Amogh Sirnoorkar Dept. of Physics and Astronomy, Purdue University, West Lafayette, IN 47907, U.S.A.    Jason W. Morphew School of Engineering Education, Purdue University, 610 Purdue Mall, West Lafayette, IN 47907, U.S.A.    Carina M. Rebello Dept. of Physics, Toronto Metropolitan University, 350 Victoria Street, Toronto, ON M5B 2K3, Canada    N. Sanjay Rebello Dept. of Physics and Astronomy / Dept. of Curriculum & Instruction, Purdue University, West Lafayette, IN 47907, U.S.A.
@@ 1 Introduction
Understanding students’ thinking in authentic STEM (science, technology, engineering, and mathematics) by integrating multiple disciplines [1, 2, 3] is essential for enhancing educational practices and outcomes. Investigating how students approach engineering design (ED)-based tasks [4, 5, 6, 7, 8] provides valuable insights into their problem-solving strategies and the integration of various disciplines [9]. Recently, the concept of Ways of Thinking (WoT) [6] has emerged as an effective framework for examining interdisciplinary problem-solving in educational settings. In the first paper [6] of this multi-part series, we introduced the Ways of Thinking in Engineering Design-based Physics (WoT4EDP) framework (see Figure 1), which integrates five key elements—design, science, mathematics, metacognitive reflection, and computational thinking—tailored specifically for introductory physics laboratory courses based on engineering design. While this paper builds on the first, we provide sufficient background and context to ensure it stands largely on its own. Table 1 provides additional details about the WoT4EDP framework.
@@
While the first paper [6] was theoretical, this paper is empirical. Our goal here is to demonstrate the application of the WoT4EDP framework to data from 14 student groups’ discussions and written reports. These artifacts [10, p.17]) stem from a student-generated engineering design (ED) problem [4] within the laboratory component of a calculus-based introductory, first semester physics course. While the first paper [6] introduced the WoT4EDP framework and highlighted its significance in undergraduate STEM education, this paper describes a qualitative study that seeks to apply the WoT4EDP framework to uncover how students engage with its five elements of the framework while tackling a student-generated engineering design challenges. By utilizing this framework, we aim to discern patterns in students’ thinking and problem-solving strategies, ultimately contributing to improved practices in STEM education. Although the five elements are interconnected, we discuss them separately in this study. While we occasionally point out instances of interconnectedness as a prelude to our next paper, a more detailed analysis will be presented in the subsequent installment of this multi-part series.
@@
Table 1: The Table describes the elements of the WoT4EDP framework as applied to our context in this study. Educators and researchers must feel empowered to modify and apply these descriptors as may fit their local contexts and pedagogical goals. More details follow in the later sections. (DBT–Design Based Thinking; SBT–Science Based Thinking; MBT–Mathematics Based thinking; MER–Metacognitive Reflection; CT - Computational Thinking.)
@@
We investigate three research questions (RQs) in this study, focusing on a student-generated engineering design (ED) problem in which students frame their own problem. The first question RQ1 examines how students articulate their problem statements. The second question RQ2 explores how design, science, mathematics, and metacognitive thinking manifest in student teams’ work. Given the emphasis on computational tasks in physics courses and the fact that computational thinking (CT) is a novel element in our WoT4EDP framework—absent in contemporary frameworks [11, 12, 13]—we dedicate the third question RQ3 to exploring CT. The research questions specific to this study are:
RQ1: What do student teams’ engineering design problem statements reveal about the aspects they consider?
RQ2: How do student teams engage with the elements of design, science, mathematics, and metacognitive reflection within the context of a student-generated engineering design problem?
RQ3: In what ways do students demonstrate computational thinking specifically related to Python coding as they tackle engineering design challenges?
While applying the WoT4EDP framework is the primary goal of this empirical study, we also take this opportunity to discuss key methodological challenges in qualitative research. Through this study, we highlight the effectiveness of qualitative methods in capturing the complexities of student thinking and problem-solving within physics education research (PER) and, more broadly, science education research [14, 15, 16]. This work, like our earlier studies [9, 6], is deeply rooted in and extends the rich tradition of qualitative research within PER [10, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. In doing so, we extend the discussion on the development and application of WoT frameworks [12, 34, 11, 6].
We now briefly outline the elements of the WoT4EDP framework in Section 2.
2 The WoT4EDP Framework
The WoT4EDP framework – Ways of Thinking for Engineering Design Problems – (see Figure 1), introduced in the first paper [6] of this multi-part series, provides a structured perspective for analyzing the diversity of student thinking approaches when addressing engineering design-based physics problems. To ensure continuity and provide necessary background, we briefly summarize the framework here before applying it in this study. The reader may refer to the first part of this series for a more detailed discussion.
The framework consists of five interconnected dimensions [6]:
  1. 1.
    Design-Based Thinking (DBT)—often referred to simply as design thinking,
  2. 2.
    Science-Based Thinking (SBT)—often referred to simply as science thinking or scientific thinking,
  3. 3.
    Mathematics-Based Thinking (MBT)—often referred to simply as mathematical thinking,
  4. 4.
    Metacognitive Reflection (MER), and
  5. 5.
    Computational Thinking (CT).
These dimensions collectively offer a comprehensive lens for examining how students integrate creativity, scientific principles, mathematical reasoning, reflective practices, and computational skills in their approaches to problem-solving. For readers’ reference, we briefly describe these dimensions and explain why each dimension is relevant to this study in Table 1.
It is important to recognize that these five elements do not always function in isolation. Students’ thinking is inherently fluid and multidimensional, making it difficult to categorize neatly into discrete silos [9]. The overlap among these dimensions will be evident as students frequently draw on multiple elements simultaneously. For example, they may use Python to model a physics concept, incorporating mathematical tools, while refining their design iteratively based on scientific reasoning. This interconnectedness warrants a deeper analysis, which is why we expanded this study into a multi-part series rather than the originally planned two-part series [6].
Despite this overlap, educators can still benefit from focusing on each aspect individually to help students develop a deeper awareness of how different ways of thinking contribute to problem-solving. This intentional emphasis strengthens students’ ability to navigate and integrate various approaches, ultimately equipping them with the skills necessary to tackle complex, real-world problems.
@@ 2 The WoT4EDP Framework
Before proceeding, we note that there is considerable debate on the definition of engineering design. The literature is replete with terms such as ‘ill-structured’, ‘ill-defined’, or even ‘wicked’ problems [6, 9]. For further discussion, we refer the reader to an insightful article by Grubbs and Strimel [35]. In this study, we consider an engineering design problem to be a real-world or practical challenge that engineers might encounter. Our students selected problems where physics principles—momentum, energy, and angular momentum—could be applied. The WoT4EDP framework is flexible enough to accommodate this perspective. Additionally, since students framed their own problems, rigid expectations could be counterproductive. Instead, we adopt a pragmatic approach for the purposes of this study: any problem students intuitively identify as an engineering design problem that integrates physics principles, mathematics, and Python coding.
@@ 3 Literature Review
The first paper [6] of this series focused on the theoretical foundations of our WoT4EDP framework, with the literature review tailored accordingly. In this second, empirical paper, we situate our work within broader literature encompassing authentic learning experiences, WoT frameworks, physics education research and qualitative research methodologies. In this section we first examine the role of authentic learning experiences, within which we position our approach of ED-based physics laboratory tasks (Section 3.1). Next, we explore the potential of frameworks such as WoT4EDP, introduced in the first paper [6], in supporting student thinking (Section 3.2). Finally, we position our work within ongoing efforts to advance the application of qualitative research methodologies in PER, as discussed in our study (Section 3.3).
@@ 3.2 Authentic Learning and WoT4EDP Framework
Montgomery et al. emphasize that to become “expert problem solvers, students must have opportunities to deliberately practice” [46, p.1] problem-solving skills. Our ED-based physics projects align with this perspective, as students engage in deliberate practice by collaborating in groups, fostering teamwork, and simulating real-world scientific and engineering practices (see Table [46, p.2]). Students tackle authentic problems that require applying fundamental physics principles, such as momentum, energy, and angular momentum [47, 9]. Students use a variety of tools—from writing equations and performing calculations by hand to automating calculations using Python—actively engaging in problem-solving that mirrors real-world contexts. Crucially, students are empowered to choose their own problems of interest, providing them with a sense of authorship or agency [48, 49, 50] that can foster deeper engagement and ownership of their learning.
@@
WoT4EDP focuses on five interconnected Ways of ThinkingDesign-Based Thinking (DBT), Science-Based Thinking (SBT), Mathematics-Based Thinking (MBT), Metacognitive Reflection (MER), and Computational Thinking (CT) [61, 62, 63, 64, 65, 66, 67, 68, 9, 47, 69, 70, 71, 72, 73, 74, 75, 76, 77]. Importantly, we acknowledge that WoT frameworks are inherently context-dependent. Even within the same instructional setting, educators may emphasize different aspects based on local constraints, instructional goals, and evolving student needs. This flexibility, which may sometimes be critiqued, is essential in respecting the complexity of educational environments, which are subject to continuous change.
@@ 3.3 Qualitative Methodology in PER - A Rich Tradition
The qualitative research community within PER produced a spirited response to the above call through articles that delve into methodological aspects. In this context, McPadden et al. [30] notably include their research team’s positionality as a validity measure, a practice echoed by Barth-Cohen et al. [21], who advocate for explicit positionality statements (which we incorporate as a reflexivity statement in our study). They emphasize the importance of transparency in ensuring rigor, a stance also reflected in Guisasola et al. [27]. Similarly, Watson and Thomas delve into aspects of  credibility,  rigor, and  quality in qualitative research. Looking ahead, we found Tschisgale et al. ’s discussion on integrating artificial intelligence into qualitative analysis methodologically relevant, as it addresses not only validity but also  scalability—an aspect we aim to explore in our future work [28].
@@ 4.1 Context
Table 2: Table shows the schedule of labs in Spring 2023. Data for this study comes from a lab section with 14 teams, submitted in Week 14. Also see Table 3. (MP – Momentum principle; EP – Energy Principle; AMP – Angular Momentum Principle)
@@
Python-based computational tasks were incorporated into the physics laboratory in a phased manner, starting in week 01. Students created their lab reports using Jupyter Notebooks [82] on Google Colab [83]. These Notebooks, accessible at no cost to anyone with a Google account, combine executable code and rich text, which facilitates running Python scripts, typing equations in LaTeX, uploading images and CSV files, and conducting data analysis without downloads or installations. Students did not need prior programming experience, as the computational tasks focused on applying physics concepts by modifying existing Python code related to the laboratory’s physical principles. Carefully designed scaffolds supported students in learning programming in the context of their experiments, which included hands-on inquiry-based activities using PASCO equipment [84] and software sensors, with data collection and analysis performed using PASCO Capstone™ software [85].

### v2

@@
This second paper in a multi-part series builds on the first, which introduced the Ways of Thinking for Engineering Design-based Physics (WoT4EDP) framework for STEM education. In this paper, we apply the framework to analyze student-groups’ discussion transcripts and written reports as they generated and solved an engineering design (ED) problem in an introductory physics laboratory. We present our findings through a detailed qualitative analysis, drawing extensively from the qualitative methods literature. We outline our analytical approach, present coding charts, and employ qualitative methods such as thematic analysis and thick description to convey our findings. Furthermore, in line with the recent recommendation in qualitative research literature, we include a reflexivity statement. In doing so, we contribute to ongoing efforts to enhance the rigor in qualitative physics education research. We examine: (i) the design and science aspects students address in their problem statements; (ii) how they engage in design-, science-, and mathematics-based thinking, as well as their metacognitive reflection while developing solutions; and (iii) how they incorporate computational thinking. Our results highlight the need for: (i) increased guidance for iterative problem framing; (ii) structured support for assessing design limitations, engaging in a feasibility study, adopting a systematic approach to applying physics and mathematics in their iterations, and engaging in metacognitive reflection; and (iii) integration of computation-based activities into laboratory tasks with appropriate scaffolding. Based on our analysis, we provide valuable insights for educators and researchers in designing physics-based engineering design tasks and promoting interdisciplinary problem-solving in STEM education.
@@
Ravishankar Chatta Subramaniam Department of Physics and Astronomy, Purdue University, West Lafayette, IN 47907, U.S.A.    Nikhil Borse Department of Physics and Astronomy, Purdue University, West Lafayette, IN 47907, U.S.A.    Winter Allen Department of Physics and Astronomy, Purdue University, West Lafayette, IN 47907, U.S.A.    Amogh Sirnoorkar Center for Advancing the Teaching and Learning of STEM, Purdue University, West Lafayette, IN 47907, U.S.A.    Jason W. Morphew School of Engineering Education, Purdue University, 610 Purdue Mall, West Lafayette, IN 47907, U.S.A.    Carina M. Rebello Department of Physics, Toronto Metropolitan University, 350 Victoria Street, Toronto, ON M5B 2K3, Canada    N. Sanjay Rebello Department of Physics and Astronomy, and Department of Curriculum and Instruction, Purdue University, West Lafayette, IN 47907, U.S.A.
@@ 1 Introduction
Understanding students’ thinking in authentic science, technology, engineering, and mathematics (STEM) learning environments by integrating multiple disciplines is essential for enhancing educational practices and outcomes [1, 2, 3]. Investigating how students approach engineering design (ED)-based tasks [4, 5, 6, 7, 8] provides valuable insights into their problem-solving strategies and the integration of various disciplines [9]. Recently, the concept of Ways of Thinking (WoT) [6] has emerged as an effective framework for examining interdisciplinary problem-solving in educational settings.
In the first paper [6] of this multi-part series, we introduced the Ways of Thinking in Engineering Design-based Physics (WoT4EDP) framework (see Figure 1, and refer Section 2 for details) which integrates five key elements—design, science, mathematics, metacognitive reflection, and computational thinking—tailored specifically for introductory physics laboratory courses based on engineering design. While this paper builds on the first, we provide sufficient background and context to ensure it stands largely on its own.
@@
While the first paper [6] was theoretical, this paper is empirical. Our goal here is to demonstrate the application of the WoT4EDP framework to data from student groups’ discussions and written reports. These artifacts [10, p.17]) stem from a student-generated engineering design (ED) problem [4, p.140] within the laboratory component of a calculus-based introductory, first semester physics course. While the first paper [6] introduced the WoT4EDP framework and highlighted its significance in undergraduate STEM education, this paper describes a qualitative approach to apply the WoT4EDP framework to uncover how students engage with its five elements while tackling student-generated engineering design challenges. By utilizing this framework, we aim to discern patterns in students’ thinking and problem-solving strategies, contributing to improved practices in STEM education. Although the five elements are interconnected, we discuss them separately in this study. While we occasionally point out instances of interconnectedness as a prelude to our next paper, a more detailed analysis will be presented in the subsequent installment of this multi-part series.
We investigate the following research questions focusing on student-generated engineering design (ED) problems.
  1. RQ1: What do student teams’ engineering design problem statements reveal about the design and science aspects they consider?
  2. RQ2: How do student teams engage with the elements of design, science, mathematics, and metacognitive reflection within the context of a student-generated engineering design problem?
  3. RQ3: In what ways do students demonstrate computational thinking specifically related to Python coding as they tackle engineering design challenges?
The first question RQ1 examines how students articulate their problem statements. The second question RQ2 explores how design, science, mathematics, and metacognitive thinking manifest in student teams’ work. Given the emphasis on computational tasks in physics courses and the fact that computational thinking (CT) is a novel element in our WoT4EDP framework—absent in contemporary frameworks on STEM ways of thinking [11, 12, 13]—we dedicate the third question RQ3 to exploring CT.
While applying the WoT4EDP framework is the primary goal of this empirical study, we also take this opportunity to discuss key methodological challenges in qualitative research. Through this study, we demonstrate the value of qualitative methods in capturing the complexities of student thinking and problem-solving within physics education research (PER) [14, 15, 16]. This work, like our earlier studies [9, 6], is deeply rooted in and extends the rich tradition of qualitative research within PER [10, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. In doing so, we extend the discussion on the development and application of WoT frameworks [12, 34, 11, 6].
This manuscript is structured as follows: In Section 2, we briefly outline the elements of the WoT4EDP framework. Section 3 reviews the background literature on authentic learning experiences, their connections to WoT4EDP, and qualitative research in PER, not only situating our study but also grounding its methodological approach. We then present the study’s context, along with methods of data collection and analysis, in Section 4. Section 5 presents the findings and discussion, where we address our research questions. In Section 6, we identify the study’s limitations. We summarize our findings in Section 8, outline implications in Section 9, and discuss future work in Section 10. While we discuss methodology throughout the paper, Section 7 is dedicated to a reflexivity statement in line with recent trends. The coding sheets in the appendixes reinforce our commitment to transparency.
2 The WoT4EDP Framework
The Ways of Thinking for Engineering Design Problems WoT4EDP framework (Figure 1), introduced in the first paper [6] of this multi-part series, provides a structured perspective for analyzing the diversity of student thinking approaches when addressing engineering design-based physics problems. To ensure continuity and provide necessary background, we briefly summarize the framework here before applying it in this study. The framework consists of the following five interconnected elements:
  1. 1.
    Design-Based Thinking (DBT) (also referred as design thinking).
  2. 2.
    Science-Based Thinking (SBT) (also referred as science thinking or scientific thinking).
  3. 3.
    Mathematics-Based Thinking (MBT) (also referred as mathematical thinking).
  4. 4.
    Metacognitive Reflection (MER), and
  5. 5.
    Computational Thinking (CT).
These dimensions collectively offer a comprehensive lens for examining how students integrate creativity, scientific principles, mathematical reasoning, reflective practices, and computational skills in their approaches to problem-solving. We briefly describe these elements and explain why each element is relevant to this study in Table 1.
@@ 2 The WoT4EDP Framework
Table 1: Elements of the WoT4EDP framework as applied to our context in this study. Educators and researchers must feel empowered to modify and apply these descriptors as may fit their local contexts and pedagogical goals. More details follow in the later sections. DBT–Design Based Thinking; SBT–Science Based Thinking; MBT–Mathematics Based thinking; MER–Metacognitive Reflection; CT - Computational Thinking.
@@
It is important to recognize that these five elements do not always function in isolation. Students’ thinking is inherently fluid and multidimensional, making it difficult to categorize neatly into discrete silos [9]. The overlap among these elements will be evident as students frequently draw on multiple elements simultaneously. For example, they may use Python to model a physics concept, incorporating mathematical tools, while refining their design iteratively based on scientific reasoning. This interconnectedness warrants a deeper analysis, which is why we expanded this study into a multi-part series rather than the originally planned two-part series [6].
Despite this overlap, educators can still benefit from focusing on each element individually to help students develop a deeper awareness of how different ways of thinking contribute to problem-solving. This intentional emphasis strengthens students’ ability to navigate and integrate various approaches, ultimately equipping them with the skills necessary to tackle complex, real-world problems.
@@
Before proceeding, we note that there is considerable debate on the definition of engineering design. The literature is replete with terms such as ‘ill-structured’, ‘ill-defined’, or even ‘wicked’ problems [6, 9]. For further discussion, we refer the reader to an insightful article by Grubbs and Strimel [35]. In this study, we define an engineering design problem as a real-world challenge engineers might face, though professional tasks are typically more complex. The problems discussed here serve as simulations for first-year engineering students. Our students selected problems where physics principles—momentum, energy, and angular momentum—apply, aligning with the flexible WoT4EDP framework. Since students framed their own problems, rigid expectations could frustrate them and detract from the task’s overall objectives. Thus, we take a pragmatic approach: any student-identified engineering design problem that integrates physics, mathematics, and Python coding is considered sufficient.
@@ 3 Literature Review
The first paper [6] of this series focused on the theoretical foundations of our WoT4EDP framework, with the literature review tailored accordingly. In this second, empirical paper, we situate our work within broader literature encompassing authentic learning experiences, WoT frameworks, physics education research, and qualitative research methodologies. In this section, we first examine the role of authentic learning experiences, within which we position our approach of ED-based physics laboratory tasks (Section 3.1). Next, we explore the potential of frameworks such as WoT4EDP, introduced in the first paper [6], in supporting student thinking in ED-based tasks (Section 3.2). Finally, we position our work within ongoing efforts to advance the application of qualitative research methodologies in PER (Section 3.3).
@@ 3.2 Authentic Learning and WoT4EDP Framework
Montgomery et al. emphasize that to become “expert problem solvers, students must have opportunities to deliberately practice” [46, p.1] problem-solving skills. Our ED-based physics projects align with this perspective: students have the opportunity to develop teamwork skills through group collaboration and engage with real-world scientific and engineering practices (see Table [46, p.2]). Students tackle authentic problems that require applying fundamental physics principles, such as momentum, energy, and angular momentum [47, 9]. Students use a variety of tools—from writing equations and performing calculations by hand to automating calculations using Python—actively engaging in problem-solving that mirrors real-world contexts. Crucially, students are empowered to choose their own problems of interest, providing them with a sense of authorship or agency [48, 49, 50] that can foster deeper engagement and ownership of their learning.
@@
WoT4EDP focuses on five interconnected Ways of ThinkingDesign-Based Thinking (DBT), Science-Based Thinking (SBT), Mathematics-Based Thinking (MBT), Metacognitive Reflection (MER), and Computational Thinking (CT) [61, 62, 63, 64, 65, 66, 67, 68, 9, 47, 69, 70, 71, 72, 73, 74, 75, 76, 77]. Importantly, we acknowledge that WoT frameworks are inherently context-dependent. Even within similar instructional setting, educators may emphasize different aspects based on local constraints, instructional goals, and evolving student needs. This flexibility, which may sometimes be critiqued, is essential in respecting the complexity of educational environments, which are subject to continuous change.
@@ 3.3 Qualitative Methodology in PER - A Rich Tradition
The qualitative research community within PER produced a spirited response to the above call through articles that delve into methodological aspects. In this context, McPadden et al. [30] notably include their research team’s positionality as a validity measure, a practice echoed by Barth-Cohen et al. [21], who advocate for explicit positionality statements (see Section 7). They emphasize the importance of transparency in ensuring rigor, a stance also reflected in Guisasola et al. [27]. Similarly, Watson and Thomas delve into aspects of  credibility,  rigor, and  quality in qualitative research. Looking ahead, we found Tschisgale et al. ’s discussion on integrating artificial intelligence into qualitative analysis methodologically relevant, as it addresses not only validity but also  scalability—an aspect we aim to explore in our future work [28].
@@ 4.1 Context
Table 2: Schedule of laboratory sessions in Spring 2023. Data for this study comes from a lab section with 14 teams, submitted in Week 14. Also see Table 3. (MP – Momentum principle; EP – Energy Principle; AMP – Angular Momentum Principle)
@@
Python-based computational tasks were incorporated into the physics laboratory in a phased manner, starting in week 01. Students created their lab reports using Jupyter Notebooks [82] in Google Colaboratory [83]. These Notebooks, accessible at no cost with a Google account, combine executable code and rich text, which facilitates running Python scripts, typing equations in LaTeX, uploading images and CSV files, and conducting data analysis without downloads or installations. Students did not need prior programming experience, as the computational tasks focused on applying physics concepts by modifying existing Python code related to the laboratory’s physical principles. Carefully designed scaffolds supported students in learning programming in the context of their experiments, which included hands-on inquiry-based activities using PASCO equipment [84] and software sensors, with data collection and analysis performed using PASCO Capstone™ software [85].